# MSVD-Indonesian: A Benchmark for Multimodal Video-Text Tasks in Indonesian

### v1

@@ 6 Discussion, Future Works and Conclusion
In this work, we constructed the MSVD-Indonesian dataset, which is the first public video-text dataset in the Indonesian language at the time of writing this paper. We conducted experiments and discussed the performance of two neural network models on our dataset for three different tasks, i.e., text-to-video retrieval, video-to-text retrieval, and video captioning. Our experimental results showed that the state-of-the-art algorithms which performed well on the English video-text dataset, i.e., the MSVD dataset, could also be applicable to our MSVD-Indonesian dataset with some modifications on the algorithms or on the parameters. Besides, incorporating a feature extractor that was pretrained on the English vision-language dataset could also help to improve the performance of the models on our Indonesian video-text dataset. Through our study, we also found that there are several potential works that can be explored for future research.

### v2

@@ 6 Discussion, Future Works and Conclusion
In this work, we constructed the MSVD-Indonesian dataset, which is the first public video-text dataset in the Indonesian language at the time of writing this paper. We conducted experiments and discussed the performance of two neural network models on our dataset for three different tasks, i.e., text-to-video retrieval, video-to-text retrieval, and video captioning. Our experimental results showed that the prior work which performed well on the English video-text dataset, i.e., the MSVD dataset, could also be applicable to our MSVD-Indonesian dataset with some modifications on the algorithms or on the parameters. Besides, incorporating a feature extractor that was pretrained on the English vision-language dataset could also help to improve the performance of the models on our Indonesian video-text dataset. Through our study, we also found that there are several potential works that can be explored for future research.